I need help solving this problem
Solve the system by back substitution
3x - 2y - 4z + 2w = 6
2y + 5z - 3w = 7
3z -4w = 0
3w = 15

Answer by umnikmo(75)

3x - 2y - 4z + 2w = 6 (1)
2y + 5z - 3w = 7 (2)
3z -4w = 0 (3)
3w = 15 (4)
from (4): w=5
from (3): z=4w/3=20/3
from (2): y=(7+3w-5z)/2=(7+3*5-5*20/3)/2=-17/3
from (1): x=(6-2w+4z+2y)/3=(6-2*5+4*20/3+2*(-17/3))/3=34/9